OpenStudy (anonymous):

A metal bar 12 cm long is rotated about one end so that the distance from the starting position of the other end is 20 cm from the ending position. Through what angle was the bar rotated?

OpenStudy (anonymous):

 \[cos\theta=\frac{A}{H}\]\[\theta=cos^{-1}\frac{A}{H}\]\[\theta=cos^{-1}\frac{12}{20}\]\[\theta=53.13^o\]

OpenStudy (vishal_kothari):

@zed r u a professor?

OpenStudy (anonymous):

tutor and student :)

OpenStudy (vishal_kothari):

which grade?

OpenStudy (anonymous):

Apparently, the answers are 112.9 degrees or 247.1 degrees. How do I get there?

OpenStudy (anonymous):

i tutor high school and am a post grad student

OpenStudy (vishal_kothari):

me in high school:)

OpenStudy (anonymous):

I think the question's like this:  How do I find x, angle of rotation?